With a strong wind behind it, a united airlines jet flies 2400 miles from los angeles to orlando in 4 hours and 45 minutes. the return trip takes 6 hours, as the plane flies into the wind. find the speed of the plane in still air, and find the wind speed to the nearest tenth mile per hour?
Answer by stanbon(75887)

With a strong wind behind it, a united airlines jet flies 2400 miles from los angeles to orlando in 4 hours and 45 minutes. the return trip takes 6 hours, as the plane flies into the wind. find the speed of the plane in still air, and find the wind speed to the nearest tenth mile per hour?
----------------
With-wind rate: 2400/(4 3/4) = 2400/(19/4) = 505.263 mph
-------
Against-wind rate: 2400/6 = 400 mph
-------
Equations:
p + w = 505.263
p - w = 400
------
Subtract and solve for "w":
2w = 105.263
---
Wind speed = 52.632 mph
==========================
Cheers,
Stan H.
============================